A solution is made by dissolving 1.25 grams of Na2SO4 in 65.0 grams of water. What is the mole fraction of solute?


1
Expert's answer
2021-10-15T02:56:21-0400

2NaOH+H2​SO4​→Na2​SO4​+2H2​O

= 65/18 = 3.61moles

125/142 = 0.88M

Total moles = 4.91

H2O = 3.61/4.91

= 0.74

Na2SO4 = 0.88/4.91

= 0.179